Home‎ > ‎Crew 165‎ > ‎

0310-DailySummaryReport

 Hello Mission Support!

 Daily Summary Report (DSR) for 10 March 2016!

SOL: 4
Person filling out Report: Ken Sullivan
Summary Title: EVA 3, water, green hab
Mission Status: All systems nominal

Sol 4 Activity Summary: Food measurement research continues. EVA 3 with NorCal Rover, research soil samples, and more fun with HAL.
  
 Anomalies in work: None
 
 Weather: Mostly sunny, light wind.   

Look ahead plan: Green Hab work tomorrow! 

 Crew Physical Status: Good!

 EVA #3: Crew members: Executive Officer Sullivan, Biologist IIzuka, GreenHab Officer Hiraki and Journalist Mori. 

Location: Around the Hab (12S-518155E-425052N, 12S-519420E-4250245N). The second coordinate was located on
 a steep hill and we did not get to that exact location. No ATV. 3 hours. Not able to get all soil samples due to time
 constraints. 

 Reports to be filed: - Commander
 Report- Engineering Report- Journalist
 Report- Daily Summary Report- EVA #3 Report
 Support Requested: No ATV Fuel.  Need Teflon Tape.